Each Book cost $11 and each toy cost $6 . how many books and toys did she buy for $157
Answer by natolino_2017(77)

In this case, we only have one equation with two variables.
Let B: number of books purchased
T: number of toys purchased.
1)11B + 6T = 157 or B = (157-6T)/11
So we make a board with the possible values of T, until the other variable (B) is also an integer.
if T = 0 then B = 157/11 (not an integer)
if T = 1 then B = 151/11 (not an integer)
if T = 2 then B = 145/11 (not an integer)
if T = 3 then B = 139/11 (not an integer)
if T = 4 then B = 133/11 (not an integer)
if T = 5 then B = 127/11 (not an integer)
if T = 6 then B = 121/11 = 11.
Here we can continue, if we need another solution (if there is another one) or we can stop the procedure.
Answer: He bought eleven books and six toys.
@natolino_